OpenStudy (anonymous):

what do these decimals have in common 5.362, 2.15, 4.794, 6.58 and what are two more decimals that can be included in this list

OpenStudy (anonymous):

When thing they have in common is that when they are rationaiized, the numerators are prime numbers. \[\{5.362,2.15,4.794,6.58\}\text{-$>$}\left\{\frac{2681}{500},\frac{43}{20},\frac{2397}{500},\frac{329}{50}\right\} \]